INVESTIGATION OF INTERNAL MOTION IN THE SYM CONFORMER OF 9-FLUORENEMETHANOL: A GAS PHASE HIGH RESOLUTION ELECTRONIC SPECTROSCOPY STUDY ^{a}

\underline{DIANE M. MITCHELL}, PHILIP J. MORGAN and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, PA 15260; JAMES A.J. FITZPATRICK, Department of Chemistry and Biomedical Engineering, Carnegie Mellon University, 4400 Fifth Avenue, Pittsburgh, PA 15217.

Last year, results were presented on the high resolution fluorescence excitation spectra of the S_{1} ← S_{0} origin transitions of two conformers of 9-fluorenemethanol (9FM). Theoretical calculations predicted two conformers, designated as symmetric and unsymmetric. Each of the two origin bands were uniquely assigned to a conformer. A splitting was observed in the origin band of the symmetric conformer, and possible explanations were presented for this splitting. One explanation involved a symmetric motion of the -OH group above the plane of the ring system, made possible by a concerted motion involving rotation about the C_{9}-C_{α} bond and rotation of oxygen about the C_{α}-O bond. Spectra have now been taken of deuterated 9FM, with the -OH group being replaced by -OD. Analysis of these results will provide further information about the observed internal motion. The sym 9FM-D origin is found to be blue-shifted relative to 9FM and displays a reduced splitting.

aWork supported by NSF (CHE-0315584)